# Consistency-based Abductive Reasoning over Perceptual Errors of Multiple Pre-trained Models in Novel Environments

### v2

@@ Introduction
Technical Preliminaries. We consider the problem of object identification for a given set of perception data $\Omega$ (Cheng and Han 2016) and assume the availability of a set $\eta$ of perception models $\mathcal{F}=\{f_{1},f_{2},\dots,f_{\eta}\}$, where each model generates a set of predictions over a shared set of $m$ object classes $\mathcal{C}=\{c_{1},...,c_{m}\}$. Under the unique name assumption [^1] See appendix for the implementation details of how we employ this assumption., every object $\omega\in\Omega$ identified by one or more models has at least one associated fact of the form $f_{i}(\omega)=c_{j}$, meaning that model $i$ has identified object $\omega$ as belonging to class $c_{j}$. We will denote the set of these facts (hereafter “observations”) as $O$.
We note that, as we are working with multiple models, there may be differences in which class is assigned a particular object. Further, some of the models may make mistakes. Hence, we introduce the predicate acceptaccept where for some model $f_{i}$ and class $c$, $accept(i,c)$ is true if we wish to accept model $f_{i}$’s results when it returns class $c$. We denote the set of all acceptance atoms with $\mathcal{H}$, and typically refer to a subset of this set as a “hypothesis”, denoted $H$. Our goal is to find a set $H$ that meets certain criteria (to be described).
@@
$error(i,c,\omega)\leftarrow(f_{i}(\omega)=c)\wedge cue(\omega)$
@@
In addition to each of the $\Pi_{i}$’s, we assume a logic program $\Pi_{helper}$ that consists of rules of the following form:
$assign(c,\omega)\leftarrow\neg error(i,c,\omega)\wedge(f_{i}(\omega)=c)\wedge accept(i,c)$
In other words, if we accept the results of a given class from a model and no error is reported for that model-class pair, then we can assign object $\omega$ class $c$. Another set of rules $\Pi_{dom}$ specifies domain knowledge. In this work, we are primarily concerned about integrity constraints that prevent a given object from being classified with conflicting classes. In this paper, rules in $\Pi_{dom}$ are of the form:
$\neg assign(c^{\prime},\omega)\leftarrow assign(c,\omega)$ (1)
We use symbol $\Pi$ to denote $\Pi_{dom}\cup\Pi_{helper}\cup(\bigcup_{i}\Pi_{i})$. Further, leveraging simple stratification and the limited use of negations of errorerror and assignassign predicates, we can implement this in a tractable monotonic logic.
@@
From a practical perspective, we may wish to allow for some (small) amount of inconsistency in the resulting output with respect to $\Pi_{dom}$. For a given hypothesis $H$, we define $Inc(H)$ as the normalized number of ground rules in $\Pi_{dom}$ not entailed by $(H\cup O\cup\Pi)\setminus\Pi_{dom}$. Given the rule format for $\Pi_{dom}$ (cf. Eq. 1), computing this is trivial (doing so in other languages is left to future work). We use the symbol $\delta\in[0,1]$ to denote an acceptable threshold for this value.
As the set of possible solutions is large, and many such solutions could be consistent with $O,\Pi\setminus\Pi_{dom}$, we employ a notion of parsimony as usual with abduction (Peng and Reggia 1990). Our parsimony function will maximize the number of atoms of the form $assign(c,\omega)$ entailed by the minimal model of $(H\cup O\cup\Pi)\setminus\Pi_{dom}$; we denote this quantity $Pred(H)$. The reason for maximization (as opposed to minimization) is threefold: $(i)$ by maximizing assigned output, we minimize the suspected errors (as the models are presumably well trained); $(ii)$ we generally seek to maximize recall; and $(iii)$ the construction of rules $\Pi_{helper}$ minimizes the impact of over-assignment of model results (as assignment occurs based not only on acceptance, but having at least one model perceive the object as belonging to the class and that it does so error-free). With this in mind, we can frame the following optimization problem:
$\max_{H\in\mathcal{H}}\ Pred(H)$
@@
$Inc(H)\leq\delta,\quad\delta\in[0,1]$
@@
$(H\cup O\cup\Pi)\setminus\Pi_{dom}\quad\text{is consistent}$
We again note that $\delta$ can be used to gauge the amount of inconsistency with domain knowledge ($\Pi_{dom}$) as opposed to inconsistency with metacognitive knowledge ($\bigcup_{i}\Pi_{i}$). There could be variance in the amount of metacognitive cues triggered by a given $\Pi_{i}$, and a user may want to vary this as well. Fortunately, the approach we used for metacognitive rule learning (Kricheli et al. 2024; Xi et al. 2024) provides an intuitive hyperparameter (denoted with $\epsilon$) that can readily be interpreted as the expected reduction in recall experienced by disregarding erroneous predictions. We examine varying $\epsilon$ as well as using a heuristic to set it automatically.
Integer Program (IP) Formulation. To solve our associated optimization, we provide an exact integer programming solution and heuristic algorithm. We first review the integer program. The goal of the IP is to find an optimal hypothesis HH—represented by a set of binary decision variables—that maximizes the total number of entailed assignments ($Pred(H)$) while ensuring the inconsistencies with our domain knowledge ($Inc(H)$) remain below the threshold $\delta$. Formally, this can be expressed as
@@
where $A_{c,\omega}$ is a binary variable indicating whether object $\omega$ is assigned to class $c$, $Con_{\omega,(c,c^{\prime})}$ is a binary variable indicating a conflict between assignments $c$ and $c^{\prime}$ for object $\omega$.
We further define the following constants and variables: $\mathit{pred}_{f,c,\omega}$ is a constant set to 1 if $(\omega,c)\in\textit{assigns}$ and $\mathit{pred}_{c}^{f}(\omega)\in\Gamma^{*}(\Pi)$, 0 otherwise; variable $X_{\omega,f,c}\in\{0,1\}$ indicates whether object $\omega$ is considered for model $f$ and class $c$; and variable $\textit{Elim}_{f,c}\in\{0,1\}$ indicates whether predictions from model $f$ for class $c$ are excluded; this last variable directly implements our choice of hypothesis $H$. Setting $\textit{Elim}_{f,c}=0$ is equivalent to including the atom $accept(f,c)$ in our hypothesis $H$, thereby trusting model $f$ for class $c$. Conversely, setting $\textit{Elim}_{f,c}=1$ excludes it. We can now present the set of constraints. First, for each $f,c,\omega$ we have constraints of the form:
@@
The IP formulation described by Equations (3)–(8) translates into a model with a number of variables and constraints dependent on the number of unique objects ($N$), models ($\mathcal{|F|}$), classes ($\mathcal{|C|}$), and integrity constraints ($|IC|$). Specifically, the IP model involves decision variables for assignments ($A_{c,\omega}$), conflict indicators ($Con_{\omega,(c,c^{\prime})}$), model-class eliminations ($Elim_{f,c}$), and object consideration ($X_{\omega,f,c}$). The total number of variables is primarily driven by the $N\times\mathcal{F}\times\mathcal{C}$ term (associated with $X_{\omega,f,c}$), resulting in an overall count in $O(N\cdot|\mathcal{F}|\cdot|\mathcal{C}|)$. Similarly, the number of constraints also scales in $O(N\cdot|\mathcal{F}|\cdot|\mathcal{C}|)$.
@@
From a knowledge representation and reasoning perspective, while solving such IP instances is NP-hard in the worst case, the specific structure of our consistency-based abduction problem often lends itself to relatively efficient resolution in practice. Our IP formulation is characterized by binary decision variables, a linear objective function (maximizing valid assignments), and a set of linear constraints. Many of these constraints exhibit a degree of locality (e.g., defining conflicts $Con_{\omega,(c,c^{\prime})}$ based on assignments $A_{c,\omega}$ for the same object $\omega$, or linking model-class considerations $X_{\omega,f,c}$ to overall assignments). This structured nature, which demonstrated efficient performance within our specific experimental configuration, often allows for practical solutions to be found within reasonable timeframes for problems of the scale explored in our experiments.
@@
Heuristic Search (HS). Our Heuristic Search (HS) approach is detailed in Algorithm 1. Given the set of all raw model predictions $P_{\textit{raw}}$, an inconsistency threshold $\delta$, and a set of EDR $\epsilon$ values to evaluate, the algorithm greedily builds a hypothesis $H$ (represented in the algorithm as the set of final predictions $S_{\textit{final}}$) by iterating through model-class pairs $(f,c)$. For each pair, it evaluates all $\epsilon$, generating a filtered prediction set $P_{\textit{new}}$ (via an implicit $\textit{GetFilteredPreds}(f,c,\epsilon,P_{\textit{raw}})$ function). It selects the $P_{\textit{new}}$ that, when added to the current solution $S_{\textit{final}}$, maximizes the size of the resulting candidate set $S_{\textit{candidate}}=S_{\textit{final}}\cup P_{\textit{new}}$, while ensuring that $\textit{ComputeInconsistency}(S_{\textit{candidate}})\leq\delta$. This chosen $P_{\textit{new}}$ for the current $(f,c)$ pair is then added to $S_{\textit{final}}$; this is analogous to deciding whether to add the atom $accept(f,c)$ to the hypothesis, based on whether this addition maximizes the number of final assignments without violating the inconsistency threshold $\delta$. The HS algorithm has a running time in $O(|\mathcal{F}|\cdot|\mathcal{C}|\cdot|E_{set}|)$, where $|\mathcal{F}|$ and $|\mathcal{C}|$ are the numbers of models and classes, and $|E_{set}|$ is the number of evaluated $\epsilon$ values—that is, the cost is polynomial with respect to these key parameters of the input. This structured, greedy method efficiently selects predictions while managing inconsistencies, rendering it suitable for large-scale problem instances.
@@
4:  $E_{set}$ (Set of EDR $\epsilon$ thresholds to evaluate)
@@
Tie-Breaker (TB) Mechanism. To ensure a single, deterministic class assignment per object from the solutions generated by our IP and HS approaches, and to resolve ambiguities where multiple class labels might be deemed equally valid post-abduction, we employ a Tie-Breaker (TB) heuristic. This TB mechanism is applied as a refinement step: if an object $\omega$ has multiple permissible class labels remaining, the TB heuristic selects the single assignments $(\omega,c)$ for which the original perception model that proposed this label reported the highest confidence score. Applying this refinement to the outputs of our core algorithms results in the IP+TB and HS+TB variants evaluated in our experiments.
@@
For evaluating the proposed approaches, we use an extended version of the Multiple Distribution Shift – Aerial (MDS-A) dataset (Ngu et al. 2025). The MDS-A dataset was generated using the AirSim (Shah et al. 2017) simulator and designed to analyze the impact of distributional shifts (caused by varying weather conditions) on object detection models in aerial imagery. The dataset consists of images captured in random positions under various weather conditions within a city environment alongside bounding boxes with one of four categories: pedestrians, vehicles, nature, and construction assigned to each bounding box. MDS-A contains six training sets, each simulating a specific weather condition: rain, snow, fog, maple leaves, dust, and a baseline (no weather effects), and three test sets containing a complex mix of weather conditions—Figure 2 illustrates an example of how the intensity distributions of weather effects vary.
@@
Both the integer programming and heuristic approaches were evaluated using a range of inconsistency thresholds: $\delta\in[0.01,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0]$. In the case of the heuristic method, the trade-off parameter $\epsilon$ was also varied over the same range. Initially, all model-class pairs were included in the optimization process, allowing the heuristic approach to adaptively determine which combinations most effectively improve prediction performance while satisfying the inconsistency constraint. The integer programming formulation followed a similar configuration, incorporating $\delta$ as a hard constraint to limit the allowable inconsistencies. Each approach was run 50 times on each test sets.
@@
All experiments were conducted on a high-performance computing system, leveraging its advanced computational capabilities. Specifically, we used two configurations: (1) high-memory node: a Dell PowerEdge R6525 equipped with AMD EPYC 7713 64-Core processors and 2TB of RAM, and (2) GPU node: a Dell PowerEdge R7525 with AMD EPYC 7413 24-Core processors, 512GB of RAM, and three NVIDIA A30 GPUs. The logical deduction process, which applies the learned EDR to raw model outputs to identify predictions and errors, was implemented using PyReason. Subsequently, the IP solutions were implemented using the PuLP library for optimization. The HS approach was implemented in Python, using libraries such as NumPy for efficient computation.
@@
As observable in Figure 4, while there is a general trend of decreasing F1-scores for all methods as the average environmental intensity increases—indicating the inherent difficulty of operating in more severe novel environments—our IP+TB approach (represented by red diamonds) consistently achieves the highest F1-scores across the entire spectrum of intensities. In nearly all instances, IP+TB surpasses the Best Individual Model, and significantly outperforms the Average Models, Majority Vote, and our Heuristic Search (HS+TB) approach in terms of F1-score. This consistent superiority, irrespective of the environmental intensity level, underscores the effectiveness of IP+TB in robustly integrating model predictions and managing inconsistencies. For instance, even at higher intensity levels where all methods experience performance degradation, IP+TB maintains a clear advantage. The same analysis for Accuracy shows comparable trends, and is provided in the supplementary material.
@@
As expected, increasing $\epsilon$ generally leads to higher precision but lower recall, demonstrating the inherent trade-off controlled by this parameter. Notably, this stricter filtering also tends to reduce the baseline level of inconsistency among the remaining predictions. This analysis highlights how tuning $\epsilon$ shapes the pool of candidate predictions subsequently processed by our IP+TB and HS+TB abduction methods aiming to maximize valid assignments under the global inconsistency constraint $\delta$. Detailed sensitivity plots for all test sets are provided in the supplementary material.
Hyperparameter Sensitivty. We analyzed the sensitivity of our approaches to the maximum inconsistency threshold hyperparameter $\delta$, which applies to IP+TB and HS+TB, and the internal recall parameter $\epsilon$ used within the IP+TB optimization. Figure 5 (right) illustrates this sensitivity analysis for the MDS-A_1 test set. The 3D plot of IP+TB show that the F1 score and accuracy reach their maximum value in the ranges of $\delta$ between 0.1 and 0.3, and for $\epsilon$ in 0.1 and 0.5. Similarly, the performance of HS+TB (center) stabilizes quickly as $\delta$ increases above a small initial value (e.g., 0.2), which is associated with the maximum level of inconsistency present in the initial test set (Figure 5, left). Detailed sensitivity plots for all test sets are provided in the appendix.
Running Time Analysis and Complexity. A key practical difference between our approaches is computational cost. Our empirical results, illustrated in a running time plot available in the supplementary material, align with the theoretical complexities of the methods. As expected, our polynomial-time HS+TB approach is significantly faster than the exact IP+TB method. While the IP solver’s running time is higher due to its exact nature, it remained tractable for the instances explored here. We observed that the average processing time per object for IP+TB did not increase steeply with more objects. We hypothesize that this is due to a combination of amortized solver overheads and structural dataset characteristics that allow for faster resolution. A more detailed study of these issues is a valuable direction for future work.

### v4

@@ Introduction
Technical Preliminaries. We consider the problem of object identification for a given set of perception data $\Omega$ (Cheng and Han 2016) and assume the availability of a set $\eta$ of perception models $\mathcal{F}=\{f_{1},f_{2},\dots,f_{\eta}\}$, where each model generates a set of predictions over a shared set of $m$ object classes $\mathcal{C}=\{c_{1},...,c_{m}\}$. Under the unique name assumption [^1] See Section 1 of the appendix for the implementation details of how we employ this assumption., every object $\omega\in\Omega$ identified by one or more models has at least one associated fact of the form $f_{i}(\omega)=c_{j}$, meaning that model $i$ has identified object $\omega$ as belonging to class $c_{j}$. We will denote the set of these facts (hereafter “observations”) as $O$.
We note that, as we are working with multiple models, there may be differences in which class is assigned a particular object. Further, some of the models may make mistakes. Hence, we introduce the predicate accept where for some model $f_{i}$ and class $c$, $\textit{accept}(i,c)$ is true if we wish to accept model $f_{i}$’s results when it returns class $c$. We denote the set of all acceptance atoms with $\mathcal{H}$, and typically refer to a subset of this set as a “hypothesis”, denoted $H$. Our goal is to find a set $H$ that meets certain criteria (to be described).
@@
$\textit{error}(i,c,\omega)\leftarrow(f_{i}(\omega)=c)\wedge\textit{cue}(\omega)$
@@
In addition to each of the $\Pi_{i}$’s, we assume a logic program $\Pi_{\textit{helper}}$ that consists of rules of the following form:
$\textit{assign}(c,\omega)\leftarrow\neg\textit{error}(i,c,\omega)\wedge(f_{i}(\omega)=c)\wedge\textit{accept}(i,c)$
In other words, if we accept the results of a given class from a model and no error is reported for that model-class pair, then we can assign object $\omega$ class $c$. Another set of rules $\Pi_{\textit{dom}}$ specifies domain knowledge. In this work, we are primarily concerned about integrity constraints that prevent a given object from being classified with conflicting classes. In this paper, rules in $\Pi_{\textit{dom}}$ are of the form:
$\neg\textit{assign}(c^{\prime},\omega)\leftarrow\textit{assign}(c,\omega)$ (1)
We use symbol $\Pi$ to denote $\Pi_{\textit{dom}}\cup\Pi_{\textit{helper}}\cup(\bigcup_{i}\Pi_{i})$. Further, leveraging simple stratification and the limited use of negations of error and assign predicates, we can implement this in a tractable monotonic logic.
@@
From a practical perspective, we may wish to allow for some (small) amount of inconsistency in the resulting output with respect to $\Pi_{\textit{dom}}$. For a given hypothesis $H$, we define $\textit{Inc}(H)$ as the normalized number of ground rules in $\Pi_{\textit{dom}}$ not entailed by $(H\cup O\cup\Pi)\setminus\Pi_{\textit{dom}}$. Given the rule format for $\Pi_{\textit{dom}}$ (cf. Eq. 1), computing this is trivial (doing so in other languages is left to future work). We use the symbol $\delta\in[0,1]$ to denote an acceptable threshold for this value.
As the set of possible solutions is large, and many such solutions could be consistent with $O,\Pi\setminus\Pi_{\textit{dom}}$, we employ a notion of parsimony as usual with abduction (Peng and Reggia 1990). Our parsimony function will maximize the number of atoms of the form $\textit{assign}(c,\omega)$ entailed by the minimal model of $(H\cup O\cup\Pi)\setminus\Pi_{\textit{dom}}$; we denote this quantity $\textit{Pred}(H)$. The reason for maximization (as opposed to minimization) is threefold: $(i)$ by maximizing assigned output, we minimize the suspected errors (as the models are presumably well trained); $(ii)$ we generally seek to maximize recall; and $(iii)$ the construction of rules $\Pi_{\textit{helper}}$ minimizes the impact of over-assignment of model results (as assignment occurs based not only on acceptance, but having at least one model perceive the object as belonging to the class and that it does so error-free). With this in mind, we can frame the following optimization problem:
$\max_{H\in\mathcal{H}}\ \textit{Pred}(H)$
@@
$\textit{Inc}(H)\leq\delta,\quad\delta\in[0,1]$
@@
$(H\cup O\cup\Pi)\setminus\Pi_{\textit{dom}}\quad\text{is consistent}$
We again note that $\delta$ can be used to gauge the amount of inconsistency with domain knowledge ($\Pi_{\textit{dom}}$) as opposed to inconsistency with metacognitive knowledge ($\bigcup_{i}\Pi_{i}$). There could be variance in the amount of metacognitive cues triggered by a given $\Pi_{i}$, and a user may want to vary this as well. Fortunately, the approach we used for metacognitive rule learning (Kricheli et al. 2024; Xi et al. 2024) provides an intuitive hyperparameter (denoted with $\epsilon$) that can readily be interpreted as the expected reduction in recall experienced by disregarding erroneous predictions. We examine varying $\epsilon$ as well as using a heuristic to set it automatically.
Integer Program (IP) Formulation. To solve our associated optimization, we provide an exact integer programming solution and heuristic algorithm. We first review the integer program. The goal of the IP is to find an optimal hypothesis HH—represented by a set of binary decision variables—that maximizes the total number of entailed assignments ($\textit{Pred}(H)$) while ensuring the inconsistencies with our domain knowledge ($\textit{Inc}(H)$) remain below the threshold $\delta$.
Formally, this can be expressed as
@@
where $A_{c,\omega}$ is a binary variable indicating whether object $\omega$ is assigned to class $c$, $\textit{Con}_{\omega,(c,c^{\prime})}$ is a binary variable indicating a conflict between assignments $c$ and $c^{\prime}$ for object $\omega$.
We further define the following constants and variables: $\mathit{pred}_{f,c,\omega}$ is a constant set to 1 if $(\omega,c)\in\textit{assigns}$ and $\mathit{pred}_{c}^{f}(\omega)\in\Gamma^{*}(\Pi)$, 0 otherwise; variable $X_{\omega,f,c}\in\{0,1\}$ indicates whether object $\omega$ is considered for model $f$ and class $c$; and variable $\textit{Elim}_{f,c}\in\{0,1\}$ indicates whether predictions from model $f$ for class $c$ are excluded; this last variable directly implements our choice of hypothesis $H$. Setting $\textit{Elim}_{f,c}=0$ is equivalent to including the atom $\textit{accept}(f,c)$ in our hypothesis $H$, thereby trusting model $f$ for class $c$. Conversely, setting $\textit{Elim}_{f,c}=1$ excludes it. We can now present the set of constraints. First, for each $f,c,\omega$ we have constraints of the form:
@@
The IP formulation described by Equations (3)–(8) translates into a model with a number of variables and constraints dependent on the number of unique objects ($N$), models ($\mathcal{|F|}$), classes ($\mathcal{|C|}$), and integrity constraints ($|\textit{IC}|$). Specifically, the IP model involves decision variables for assignments ($A_{c,\omega}$), conflict indicators ($\textit{Con}_{\omega,(c,c^{\prime})}$), model-class eliminations ($\textit{Elim}_{f,c}$), and object consideration ($X_{\omega,f,c}$). The total number of variables is primarily driven by the $N\times\mathcal{F}\times\mathcal{C}$ term (associated with $X_{\omega,f,c}$), resulting in an overall count in $O(N\cdot|\mathcal{F}|\cdot|\mathcal{C}|)$. Similarly, the number of constraints also scales in $O(N\cdot|\mathcal{F}|\cdot|\mathcal{C}|)$.
@@
From a knowledge representation and reasoning perspective, while solving such IP instances is NP-hard in the worst case, the specific structure of our consistency-based abduction problem often lends itself to relatively efficient resolution in practice. Our IP formulation is characterized by binary decision variables, a linear objective function (maximizing valid assignments), and a set of linear constraints. Many of these constraints exhibit a degree of locality (e.g., defining conflicts $\textit{Con}_{\omega,(c,c^{\prime})}$ based on assignments $A_{c,\omega}$ for the same object $\omega$, or linking model-class considerations $X_{\omega,f,c}$ to overall assignments). This structured nature, which demonstrated efficient performance within our specific experimental configuration, often allows for practical solutions to be found within reasonable timeframes for problems of the scale explored in our experiments.
@@
Heuristic Search (HS). Our Heuristic Search (HS) approach is detailed in Algorithm 1. Given the set of all raw model predictions $P_{\textit{raw}}$, an inconsistency threshold $\delta$, and a set of EDR $\epsilon$ values to evaluate, the algorithm greedily builds a hypothesis $H$ (represented in the algorithm as the set of final predictions $S_{\textit{final}}$) by iterating through model-class pairs $(f,c)$. For each pair, it evaluates all $\epsilon$, generating a filtered prediction set $P_{\textit{new}}$ (via an implicit $\textit{GetFilteredPreds}(f,c,\epsilon,P_{\textit{raw}})$ function). It selects the $P_{\textit{new}}$ that, when added to the current solution $S_{\textit{final}}$, maximizes the size of the resulting candidate set $S_{\textit{candidate}}=S_{\textit{final}}\cup P_{\textit{new}}$, while ensuring that $\textit{ComputeInconsistency}(S_{\textit{candidate}})\leq\delta$. This chosen $P_{\textit{new}}$ for the current $(f,c)$ pair is then added to $S_{\textit{final}}$; this is analogous to deciding whether to add the atom $\textit{accept}(f,c)$ to the hypothesis, based on whether this addition maximizes the number of final assignments without violating the inconsistency threshold $\delta$. The HS algorithm has a running time in $O(|\mathcal{F}|\cdot|\mathcal{C}|\cdot|E_{set}|)$, where $|\mathcal{F}|$ and $|\mathcal{C}|$ are the numbers of models and classes, and $|E_{set}|$ is the number of evaluated $\epsilon$ values—that is, the cost is polynomial with respect to these key parameters of the input. This structured, greedy method efficiently selects predictions while managing inconsistencies, rendering it suitable for large-scale problem instances.
@@
4:  $E_{\textit{set}}$ (Set of EDR $\epsilon$ thresholds to evaluate)
@@
Tie-Breaker (TB) Mechanism. To ensure deterministic class assignment per object and resolve ambiguities where multiple labels remain valid after abduction, we apply a Tie-Breaker (TB) heuristic. For any object $\omega$ with several admissible labels, TB selects the pair $(\omega,c)$ proposed by the perception model with the highest confidence. This refinement yields the IP+TB and HS+TB variants in our experiments.
@@
For evaluating the proposed approaches, we use an extended version of the Multiple Distribution Shift -– Aerial (MDS-A) dataset (Ngu et al. 2025). The MDS-A dataset was generated using the AirSim (Shah et al. 2017) simulator and designed to analyze the impact of distributional shifts (caused by varying weather conditions) on object detection models in aerial imagery. The dataset consists of images captured in random positions under various weather conditions within a city environment alongside bounding boxes with one of four categories: pedestrians, vehicles, nature, and construction assigned to each bounding box. MDS-A contains six training sets, each simulating a specific weather condition: rain, snow, fog, maple leaves, dust, and a baseline (no weather effects), and three test sets containing a complex mix of weather conditions—Figure 2 illustrates an example of how the intensity distributions of weather effects vary.
@@
Both the integer programming and heuristic approaches were evaluated using a range of inconsistency thresholds: $\delta\in[0.01,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0]$. For the heuristic method, the trade-off parameter $\epsilon$ was also varied over the same range. Initially, all model-class pairs were included in the optimization process, allowing the heuristic approach to adaptively determine which combinations most effectively improve prediction performance while satisfying the inconsistency constraint. The integer programming formulation followed a similar configuration, incorporating $\delta$ as a hard constraint to limit the allowable inconsistencies. Each approach was run 50 times on each test sets.
@@
All experiments were conducted on a high-performance computing system, leveraging its advanced computational capabilities. Specifically, we used two configurations: (1) high-memory node: a Dell PowerEdge R6525 equipped with AMD EPYC 7713 64-Core processors and 2TB of RAM, and (2) GPU node: a Dell PowerEdge R7525 with AMD EPYC 7413 24-Core processors, 512GB of RAM, and three NVIDIA A30 GPUs. The logical deduction process, which applies the learned EDR to raw model outputs to identify predictions and errors, was implemented using PyReason. Subsequently, the IP solutions were implemented using the PuLP library for optimization.
@@
As observable in Figure 4, while there is a general trend of decreasing F1-scores for all methods as the average environmental intensity increases—indicating the inherent difficulty of operating in more severe novel environments—our IP+TB approach (represented by red diamonds) consistently achieves the highest F1-scores across the entire spectrum of intensities. In nearly all instances, IP+TB surpasses the Best Individual Model, and significantly outperforms the Average Models, Majority Vote, and our Heuristic Search (HS+TB) approach in terms of F1-score. This consistent superiority, irrespective of the environmental intensity level, underscores the effectiveness of IP+TB in robustly integrating model predictions and managing inconsistencies. For instance, even at higher intensity levels where all methods experience performance degradation, IP+TB maintains a clear advantage. The same analysis for Accuracy shows comparable trends, and is provided in the appendix.
@@
As expected, increasing $\epsilon$ generally leads to higher precision but lower recall, demonstrating the inherent trade-off controlled by this parameter. Notably, this stricter filtering also tends to reduce the baseline level of inconsistency among the remaining predictions. This analysis highlights how tuning $\epsilon$ shapes the pool of candidate predictions subsequently processed by our IP+TB and HS+TB abduction methods aiming to maximize valid assignments under the global inconsistency constraint $\delta$. Detailed sensitivity plots for all test sets are provided in the appendix.
Hyperparameter Sensitivity. We analyzed the sensitivity of our approaches to the maximum inconsistency threshold hyperparameter $\delta$, which applies to IP+TB and HS+TB, and the internal recall parameter $\epsilon$ used within the IP+TB optimization. Figure 5 (right) illustrates this analysis for the MDS-A_1 test set. The 3D plot of IP+TB show that the F1 score and accuracy reach their maximum value in the ranges of $\delta$ between 0.1 and 0.3, and for $\epsilon$ in 0.1 and 0.5. Similarly, the performance of HS+TB (center) stabilizes quickly as $\delta$ increases above a small initial value (e.g., 0.2), which is associated with the maximum level of inconsistency present in the initial test set (Figure 5, left); detailed sensitivity plots for all test sets are provided in the appendix.
Running Time Analysis and Complexity. A key practical difference between our approaches is computational cost. Our empirical results, illustrated in a running time plot available in the appendix, align with the theoretical complexities of the methods. As expected, our polynomial-time HS+TB approach is significantly faster than the exact IP+TB method. While the IP solver’s running time is higher due to its exact nature, it remained tractable for the instances explored here. We observed that the average processing time per object for IP+TB did not increase steeply with more objects. We hypothesize that this is due to a combination of amortized solver overheads and structural dataset characteristics that allow for faster resolution. A detailed study of these issues is future work.